3. A 5kg block is connected to a 3kg block by means of a string that passes over a pulley. Find the (a) unbalanced force, (b) acceleration of the blocks, and (c) tension in the string. Assume that the pulley and the surfaces are frictionless.


1
Expert's answer
2022-01-10T09:07:49-0500

(a) The unbalanced force is


"F_u=g(M-m)=19.6\text{ N}."


(b) acceleration of the blocks:


"a=F/(M+m)=g\dfrac{M-m}{M+m}=2.45\text{ m/s}^2."


(c) tension in the string:


"T=m(g+a)=M(g-a)= 36.75\text{ N}."